OpenStudy (anonymous):

The length of a rectangle is 3 centimeters more than 3 times the width. If the perimeter of the rectangle is 46 centimeters, find the dimensions of the rectangle

OpenStudy (anonymous):

w=width 3w+3=length Perimeter=2l+2w Fill in the blanks

OpenStudy (shamil98):

P = 2L + 2W L = 3 + W 2(3+w) + 2W = 46 solve for width, and then find the length by inputting the value back into 2L + 2W = P

OpenStudy (shamil98):

I made an error L = 3 + 3w *

OpenStudy (anonymous):

are you sure about the width? 46=2w+2(3w+3) 46=2w+6w+6 46=8w+6 40=8w 5=w

OpenStudy (anonymous):

length =18

OpenStudy (anonymous):

let x be width than length=3x+3 2(length + width)=46 2(3x+3+x)=46 4x+3=23 x=5 width=5 length =18

OpenStudy (anonymous):

Thank you